# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### 06 Mar 2026 v4

@@ Appendix B The bosonic birthday paradox for the shallow average-case circuits
Throughout our work, we consider only collision-free outcomes in our average-case hardness arguments instead of all possible outcomes. In this appendix, we examine whether the bosonic birthday paradox [aaronson2011computational] holds for our average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$. More specifically, we investigate whether collision-free outcomes are observed with high probability in the so-called “dilute” regime, where the number of modes satisfies $M=\Omega(N^{2})$.
@@
Accordingly, to avoid this issue and claim the classical hardness result using only collision-free outcomes, the bosonic birthday paradox is crucial, which suppresses the occurrence of collision outcomes over the experiments. This property holds for global Haar random circuits in the dilute regime $M=\Omega(N^{2})$ [aaronson2011computational], but not guaranteed for any other random circuit ensemble, including our circuit ensemble. In fact, collision-free outcomes dominate the outcome space in the regime $M=\Omega(N^{2})$ since the portion of collision-free outcomes over all possible outcomes can be bounded as $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$. Accordingly, one might expect that the bosonic birthday paradox automatically holds for any circuit ensemble. However, even in this regime, the probability weight might be highly concentrated on collision outcomes for some circuit ensembles (i.e., anticoncentration is not guaranteed for arbitrary circuit ensembles), such that the probability weight of the collision-free outcomes is still too small. Hence, it is crucial to show that the bosonic birthday paradox for our circuit ensemble, such that our circuit ensemble outputs collision-free outcomes with high probability.
Based on this understanding, we now argue the bosonic birthday paradox for the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ in the dilute regime $M=\Omega(N^{2})$. We first review a lemma given in Ref. [aaronson2011computational, Lemma C.1], which is indeed a crucial ingredient for the bosonic birthday paradox.
Lemma 4 (Unitary Pigeonhole Principle [aaronson2011computational]).
@@ Appendix C Proof of Theorem 3 (worst-case hardness)
Lemma 5 (Brod [brod2015complexity], revised).
@@
In this proof, we revise the main result in [brod2015complexity], incorporating a more quantitative analysis on the coefficient $c_{Q}$, which later determines the imprecision level allowed for the worst-case hardness of shallow-depth Boson Sampling. To do so, we first go through the main result in [brod2015complexity].
Given an arbitrary poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state composed of maximally $m_{Q}=\text{poly}(n)$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, with an explicit measurement basis sequence for the universal gate set {CX, T, H} to simulate $Q$ [raussendorf2001one, broadbent2009universal, childs2005unified] (see, e.g., Fig. 4 (a)). Since the measurement outcomes for gate implementation are random, one needs adaptive measurements based on these measurement outcomes to implement $Q$ (which is indeed the essence of MBQC). However, if post-selecting desired measurement outcomes are allowed, one can rather post-select “good" outcomes that do not require any adaptive measurements to implement $Q$.
@@
where $p_{k}$ denotes post-selection probability to implement $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [knill2002quantum]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. Hence, given the post-selection probability $p_{k}$ for each gate, by counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. To obtain the size of $c_{Q}$, note that we set the photon number $N=\Theta(m_{Q})$ to implement $m_{Q}$ qubit brickwork qubit circuit. As this $m_{Q}$ qubit brickwork qubit circuit can implement at most $O(m_{Q})$ number of gates, $Q$ has at most $O(m_{Q})=O(N)$ number of gates in the universal gate set {CX, T, H}. Also, the number of CZ gates is given as $O(N)$, and thus the total number of gates to implement $Q$ is at most $\sum_{k}\Gamma_{k}=O(N)$. Therefore, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, since all the post-selection probabilities $p_{k}$ are given as constants. This concludes the proof.
Next, we use the proposition by [kondo2022quantum] which states that estimating the output probability of any BQP circuit is #P-hard for a certain additive imprecision.
Lemma 6 (Kondo et al [kondo2022quantum]).
@@
In the following, we prove Lemma 7, showing that the constant depth worst-case linear optical circuit $C_{0}$ argued in [brod2015complexity] and Lemma 5 can be encoded in the Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by properly allocating mode indices. To summarize, this can be done in two steps: First, we show how we can map the worst-case linear optical circuit $C_{0}$ into a 3-dimensional “grid" linear optical circuit (i.e., a circuit with only nearest neighbor interactions for each dimension, with open boundaries). Next, we show that a single parallel application of any grid-structured linear optical circuit can be implemented in the unit butterfly circuit architecture $\mathcal{B}$ under a mode index permutation. We then combine these arguments, showing that the worst-case circuit $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$.
@@ Appendix E About Cayley transform
As we depicted in the main text, our goal is to perturb the average-case circuit with the worst-case circuit, and it is crucial to choose a proper circuit perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [movassagh2023hardness, bouland2022noise, kondo2022quantum] for the hardness proposals of the random circuit sampling.
Definition 8 (Cayley transform [movassagh2023hardness]).
@@
We should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation described in Definition 9, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance. In fact, Ref. [movassagh2023hardness] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C}$ is small for sufficiently small perturbation $\theta$ compared to the inverse gate number of the circuit architecture $\mathcal{A}$.
Lemma 9 (Movassagh [movassagh2023hardness]).
@@ Appendix F Proof of Theorem 4 (average-case hardness)
where $D_{\rm{TV}}$ denotes total variation distance. The inequality in Eq. (33) is evident based on the two facts: We can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [bouland2019complexity], and the equal permutation on the circuit does not affect the total variation distance over the circuit space. In the right-hand side of Eq. (33), by Lemma 9, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
Therefore, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (F) is near $\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. To avoid this issue, we employ the strategy from Ref. [bouland2022noise], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (38) below) with high probability over random circuit instances.
@@
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm that requires the NP oracle, recently proposed in Ref. [bouland2022noise].
Theorem 10 (Robust Berlekamp-Welch [bouland2022noise]).
@@ Appendix G Proof of Lemma 3
for all $k>2$. Also, using Stockmeyer’s algorithm [stockmeyer1985approximation] whose complexity is in $\rm{BPP}^{\rm{NP}}$, obtaining the estimate $\tilde{p}_{\bm{s}}(C)$ of $\bar{p}_{\bm{s}}(C)$ satisfying
@@ Appendix H On average-case hardness including collision outcomes
Also, as the final imprecision level increases by the estimated value of $p_{\bm{s}_{1}}(U_{0})^{-1}$, it must be promised that $p_{\bm{s}_{1}}(U_{0})$ is lower bounded. More precisely, $p_{\bm{s}_{1}}(U_{0})$ must be at least as large as the allowed imprecision level of average-case estimation (i.e., imprecision level of the oracle $\mathcal{O}$), commonly referred to as the anti-concentration property [aaronson2011computational, bouland2019complexity, Bremner2016average, Hangleiter2018anticoncentration, Bermejo2018architectures, Dalzell2022random]. Nevertheless, we believe that $p_{\bm{s}_{1}}(U_{0})$ is at least as large as the allowed imprecision level for most of $U_{0}$, based on two reasons: First, the imprecision level required by our current average-case hardness result is sufficiently small. Also, although this has only been demonstrated in finite-sized numerical experiments, random circuits for our circuit architecture closely resemble global Haar-random unitary circuits [go2024exploring], which are strongly conjectured to exhibit the anti-concentration property [aaronson2011computational].
@@
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw $N$ balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [ross2014introduction]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We emphasize that the distribution of the classical balls-into-bins problem is not equivalent to the uniform distribution over the possible outcomes of Boson Sampling, because the classical balls are distinguishable while identical bosons are indistinguishable in nature (and indeed, bosons are even more “gregarious" than classical balls [aaronson2011computational]). Nevertheless, this example is sufficient to provide an insight that a sufficiently large number of singletons would appear in a large fraction of Boson Sampling outcomes.
We conclude this appendix by introducing an alternative approach, based on the closely related result in [bouland2023complexity] that establishes a hardness argument for general Boson Sampling in the saturated regime. Note that our convention on singletons closely parallels the notion of “clicks" introduced in [bouland2023complexity], which refers to the number of modes that are occupied by a non-zero number of photons. Analogous to our analysis, their key observation is that, even in the saturated regime, most outcomes have a sufficiently large number of clicks. Based on this observation, they show that, given an $N$-photon outcome $\bm{s}$ with collisions that has a sufficient number of clicks $N_{0}$, there exists a poly-time constructible matrix $A\in\{0,1\}^{N_{0}\times N}$ such that computing $\text{Per}(A_{\bm{s}})$ is worst-case #P-hard, for $A_{\bm{s}}$ being a submatrix of $A$ by repeating rows according to $\bm{s}$. Then they establish a worst-to-average-case reduction by perturbing $A$ into their average-case matrix and performing polynomial interpolation (for details, see [bouland2023complexity]).
@@
To show the bosonic birthday paradox property for the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, we show numerically that the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ samples sufficiently many collision-free outcomes over the experiments. Specifically, we prepare random unitary circuits drawn from the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, sample the outcomes by classical Boson Sampling algorithm for each randomly chosen circuit, and compute the ratio of the collision-free outcomes over all the sampled outcomes for each randomly chosen circuit. To evaluate whether these local random circuits output sufficiently many collision-free outcomes, as a comparison, we also prepare the global Haar random circuits, which have the bosonic birthday paradox property as depicted in [aaronson2011computational]. In other words, if the ratio of collision-free outcomes for local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ is as large as the ratio of collision-free outcomes for global Haar random circuit, this clearly gives evidence of the bosonic birthday paradox for our circuit architecture.
@@
Figure 6: The average ratio of collision-free outcomes from (a) 256256-mode and (b) 3232-mode unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. We sample 500 unitary circuits from each of the circuit ensembles, and we sample 500 outcomes on input $N=4,8,12,16$ photons for each of the randomly chosen circuits. Then we count the collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes over 500 randomly chosen circuits for each of the circuit ensembles. We present in the figure the average ratio of collision-free outcomes for local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=1,2,3$ and global Haar random circuit ensemble, for the photon number given by $N=4,8,12,16$ (the error bar represents standard deviation). For all of the repetition numbers $q=1,2,3$ and the photon numbers $N=4,8,12,16$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost equivalent average collision-free ratio to the global Haar random unitary circuits.
@@
More specifically, we consider $256$-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first $N$ modes over $256$ modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [clifford2018classical]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each $N$ and circuit ensemble, where the average is taken over the $500$ sampled unitary circuits for each circuit ensemble.

### 11 Mar 2026 v5

@@ Appendix B The bosonic birthday paradox for the shallow average-case circuits
Throughout our work, we consider only collision-free outcomes in our average-case hardness arguments instead of all possible outcomes. In this appendix, we examine whether the bosonic birthday paradox [1] holds for our average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$. More specifically, we investigate whether collision-free outcomes are observed with high probability in the so-called “dilute” regime, where the number of modes satisfies $M=\Omega(N^{2})$.
@@
Accordingly, to avoid this issue and claim the classical hardness result using only collision-free outcomes, the bosonic birthday paradox is crucial, which suppresses the occurrence of collision outcomes over the experiments. This property holds for global Haar random circuits in the dilute regime $M=\Omega(N^{2})$ [1], but not guaranteed for any other random circuit ensemble, including our circuit ensemble. In fact, collision-free outcomes dominate the outcome space in the regime $M=\Omega(N^{2})$ since the portion of collision-free outcomes over all possible outcomes can be bounded as $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$. Accordingly, one might expect that the bosonic birthday paradox automatically holds for any circuit ensemble. However, even in this regime, the probability weight might be highly concentrated on collision outcomes for some circuit ensembles (i.e., anticoncentration is not guaranteed for arbitrary circuit ensembles), such that the probability weight of the collision-free outcomes is still too small. Hence, it is crucial to show that the bosonic birthday paradox for our circuit ensemble, such that our circuit ensemble outputs collision-free outcomes with high probability.
Based on this understanding, we now argue the bosonic birthday paradox for the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ in the dilute regime $M=\Omega(N^{2})$. We first review a lemma given in Ref. [1, Lemma C.1], which is indeed a crucial ingredient for the bosonic birthday paradox.
Lemma 4 (Unitary Pigeonhole Principle [1]).
@@ Appendix C Proof of Theorem 3 (worst-case hardness)
Lemma 5 (Brod [33], revised).
@@
In this proof, we revise the main result in [33], incorporating a more quantitative analysis on the coefficient $c_{Q}$, which later determines the imprecision level allowed for the worst-case hardness of shallow-depth Boson Sampling. To do so, we first go through the main result in [33].
Given an arbitrary poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state composed of maximally $m_{Q}=\text{poly}(n)$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, with an explicit measurement basis sequence for the universal gate set {CX, T, H} to simulate $Q$ [52, 53, 54] (see, e.g., Fig. 4 (a)). Since the measurement outcomes for gate implementation are random, one needs adaptive measurements based on these measurement outcomes to implement $Q$ (which is indeed the essence of MBQC). However, if post-selecting desired measurement outcomes are allowed, one can rather post-select “good" outcomes that do not require any adaptive measurements to implement $Q$.
@@
where $p_{k}$ denotes post-selection probability to implement $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [56]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. Hence, given the post-selection probability $p_{k}$ for each gate, by counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. To obtain the size of $c_{Q}$, note that we set the photon number $N=\Theta(m_{Q})$ to implement $m_{Q}$ qubit brickwork qubit circuit. As this $m_{Q}$ qubit brickwork qubit circuit can implement at most $O(m_{Q})$ number of gates, $Q$ has at most $O(m_{Q})=O(N)$ number of gates in the universal gate set {CX, T, H}. Also, the number of CZ gates is given as $O(N)$, and thus the total number of gates to implement $Q$ is at most $\sum_{k}\Gamma_{k}=O(N)$. Therefore, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, since all the post-selection probabilities $p_{k}$ are given as constants. This concludes the proof.
Next, we use the proposition by [43] which states that estimating the output probability of any BQP circuit is #P-hard for a certain additive imprecision.
Lemma 6 (Kondo et al [43]).
@@
In the following, we prove Lemma 7, showing that the constant depth worst-case linear optical circuit $C_{0}$ argued in [33] and Lemma 5 can be encoded in the Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by properly allocating mode indices. To summarize, this can be done in two steps: First, we show how we can map the worst-case linear optical circuit $C_{0}$ into a 3-dimensional “grid" linear optical circuit (i.e., a circuit with only nearest neighbor interactions for each dimension, with open boundaries). Next, we show that a single parallel application of any grid-structured linear optical circuit can be implemented in the unit butterfly circuit architecture $\mathcal{B}$ under a mode index permutation. We then combine these arguments, showing that the worst-case circuit $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$.
@@ Appendix E About Cayley transform
As we depicted in the main text, our goal is to perturb the average-case circuit with the worst-case circuit, and it is crucial to choose a proper circuit perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [42, 28, 43] for the hardness proposals of the random circuit sampling.
Definition 8 (Cayley transform [42]).
@@
We should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation described in Definition 9, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance. In fact, Ref. [42] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C}$ is small for sufficiently small perturbation $\theta$ compared to the inverse gate number of the circuit architecture $\mathcal{A}$.
Lemma 9 (Movassagh [42]).
@@ Appendix F Proof of Theorem 4 (average-case hardness)
where $D_{\rm{TV}}$ denotes total variation distance. The inequality in Eq. (33) is evident based on the two facts: We can interpret the total variation distance as the supremum over events of the difference in probabilities of those events (Viz., circuits corresponding to the failure) [41], and the equal permutation on the circuit does not affect the total variation distance over the circuit space. In the right-hand side of Eq. (33), by Lemma 9, $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ is $O(m\theta)$. By setting $0\leq\theta\leq\Delta$ with $\Delta=O(m^{-1})$, we can upper bound $D_{\rm{TV}}(\mathcal{H}_{\mathcal{A},\theta}^{C_{p}},\mathcal{H}_{\mathcal{A}})$ by an arbitrarily small constant.
@@
Therefore, $p_{\bm{s}_{p}}(V(\theta)\bm{P}_{1})$ is very close to the degree $d=4mN$ polynomial $P(\theta)$ in $\theta\in[0,\Delta]$, which allows us to use polynomial interpolation technique for $P(\theta)$. Specifically, we obtain estimations of $P(\theta)$ for different values of $\theta\in[0,\Delta]$ by querying the oracle $\mathcal{O}$, use polynomial interpolation for given $P(\theta)$ values to estimate $P(1)$, and infer the value $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})=p_{\bm{s}_{0}}(C_{0})$ by multiplying $Q(1)^{-1}$. However, $Q(1)^{-1}$ becomes arbitrarily large for the case that even a single $\phi_{i,j}$ in Eq. (F) is near $\pm\pi$, which will arbitrarily enlarge the imprecision of the approximation of $p_{\bm{s}_{p}}(V(1)\bm{P}_{1})$. To avoid this issue, we employ the strategy from Ref. [28], which only considers the case that all $\phi_{i,j}$ values of randomly chosen gates $\{H_{i}\}_{i=1}^{m}$ are in $\pi+\zeta,\pi-\zeta]$, and regards the other case as failure. This happens with probability at least $1-O(m\zeta)$ over the random circuit instances. By setting $\zeta=O(m^{-1})$, we can make $O(m\zeta)$ arbitrarily small constant, and as a result, we can upper bound $Q(1)^{-1}$ (see Eq. (38) below) with high probability over random circuit instances.
@@
For the polynomial interpolation, we employ the Robust Berlekamp-Welch algorithm that requires the NP oracle, recently proposed in Ref. [28].
Theorem 10 (Robust Berlekamp-Welch [28]).
@@ Appendix G Proof of Lemma 3
for all $k>2$. Also, using Stockmeyer’s algorithm [48] whose complexity is in $\rm{BPP}^{\rm{NP}}$, obtaining the estimate $\tilde{p}_{\bm{s}}(C)$ of $\bar{p}_{\bm{s}}(C)$ satisfying
@@ Appendix H On average-case hardness including collision outcomes
Also, as the final imprecision level increases by the estimated value of $p_{\bm{s}_{1}}(U_{0})^{-1}$, it must be promised that $p_{\bm{s}_{1}}(U_{0})$ is lower bounded. More precisely, $p_{\bm{s}_{1}}(U_{0})$ must be at least as large as the allowed imprecision level of average-case estimation (i.e., imprecision level of the oracle $\mathcal{O}$), commonly referred to as the anti-concentration property [1, 41, 57, 58, 59, 60]. Nevertheless, we believe that $p_{\bm{s}_{1}}(U_{0})$ is at least as large as the allowed imprecision level for most of $U_{0}$, based on two reasons: First, the imprecision level required by our current average-case hardness result is sufficiently small. Also, although this has only been demonstrated in finite-sized numerical experiments, random circuits for our circuit architecture closely resemble global Haar-random unitary circuits [47], which are strongly conjectured to exhibit the anti-concentration property [1].
@@
Let us first denote the random variable $X_{i}$ as the number of balls in $i$-th bin (such that $\sum_{i=1}^{M}X_{i}=N$), when we throw $N$ balls into $M$ bins uniformly at random. Here, our goal is to verify that over the experiments, it is highly unlikely to observe the outcomes that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$. Since all the variables $X_{i}$ are correlated due to the constraint $\sum_{i=1}^{M}X_{i}=N$, instead, we consider i.i.d. random variable $Y_{i}$ each drawn from Poisson distribution $P(\frac{N}{M})$, such that $\operatorname*{\mathbb{E}}[\sum_{i=1}^{M}Y_{i}]=N$ and $\Pr[\sum_{i=1}^{M}Y_{i}=N]=e^{-N}\frac{N^{N}}{N!}=\Theta(\frac{1}{\sqrt{N}})$. The motivation for using Poissonian random variables is that the distribution of $(Y_{1},\dots,Y_{M})$ conditioned on the sum $\sum_{i=1}^{M}Y_{i}=N$ is equivalent to the distribution of the balls-into-bins problem $(X_{1},\dots,X_{M})$ (see, e.g., Ref. [61]). Accordingly, if it is exponentially unlikely that the number of $Y_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(Y_{1},\dots,Y_{M})$, this indicates that it is still exponentially unlikely that the number of $X_{i}=1$ is smaller than $cN^{\lambda}$ as we sample $(X_{1},\dots,X_{M})$, given that the post-selection probability for $\sum_{i=1}^{M}Y_{i}=N$ is at most inverse polynomial (i.e., $\Theta(\frac{1}{\sqrt{N}})$).
@@
We emphasize that the distribution of the classical balls-into-bins problem is not equivalent to the uniform distribution over the possible outcomes of Boson Sampling, because the classical balls are distinguishable while identical bosons are indistinguishable in nature (and indeed, bosons are even more “gregarious" than classical balls [1]). Nevertheless, this example is sufficient to provide an insight that a sufficiently large number of singletons would appear in a large fraction of Boson Sampling outcomes.
We conclude this appendix by introducing an alternative approach, based on the closely related result in [29] that establishes a hardness argument for general Boson Sampling in the saturated regime. Note that our convention on singletons closely parallels the notion of “clicks" introduced in [29], which refers to the number of modes that are occupied by a non-zero number of photons. Analogous to our analysis, their key observation is that, even in the saturated regime, most outcomes have a sufficiently large number of clicks. Based on this observation, they show that, given an $N$-photon outcome $\bm{s}$ with collisions that has a sufficient number of clicks $N_{0}$, there exists a poly-time constructible matrix $A\in\{0,1\}^{N_{0}\times N}$ such that computing $\text{Per}(A_{\bm{s}})$ is worst-case #P-hard, for $A_{\bm{s}}$ being a submatrix of $A$ by repeating rows according to $\bm{s}$. Then they establish a worst-to-average-case reduction by perturbing $A$ into their average-case matrix and performing polynomial interpolation (for details, see [29]).
@@
To show the bosonic birthday paradox property for the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, we show numerically that the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ samples sufficiently many collision-free outcomes over the experiments. Specifically, we prepare random unitary circuits drawn from the local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$, sample the outcomes by classical Boson Sampling algorithm for each randomly chosen circuit, and compute the ratio of the collision-free outcomes over all the sampled outcomes for each randomly chosen circuit. To evaluate whether these local random circuits output sufficiently many collision-free outcomes, as a comparison, we also prepare the global Haar random circuits, which have the bosonic birthday paradox property as depicted in [1]. In other words, if the ratio of collision-free outcomes for local random circuit ensemble over $(\mathcal{B}\mathcal{B}^{*})^{q}$ is as large as the ratio of collision-free outcomes for global Haar random circuit, this clearly gives evidence of the bosonic birthday paradox for our circuit architecture.
@@
Figure 6: The average ratio of collision-free outcomes from (a) $256$-mode and (b) $32$-mode unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. We sample 500 unitary circuits from each of the circuit ensembles, and we sample 500 outcomes on input $N=4,8,12,16$ photons for each of the randomly chosen circuits. Then we count the collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes over 500 randomly chosen circuits for each of the circuit ensembles. We present in the figure the average ratio of collision-free outcomes for local random circuits over $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=1,2,3$ and global Haar random circuit ensemble, for the photon number given by $N=4,8,12,16$ (the error bar represents standard deviation). For all of the repetition numbers $q=1,2,3$ and the photon numbers $N=4,8,12,16$, local random circuits for $(\mathcal{B}\mathcal{B}^{*})^{q}$ exhibits almost equivalent average collision-free ratio to the global Haar random unitary circuits.
@@
More specifically, we consider $256$-mode random unitary circuits drawn from local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ over the $q$-Kaleidoscope circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ with the repetition number $q=1,2,3$, and global Haar random circuit ensemble. First, we sample 500 unitary circuits from each of the circuit ensembles described above. For each of the randomly chosen circuits, on input $N=4,8,12,16$ photons in the first $N$ modes over $256$ modes, we sample 500 outcomes using the Clifford-Clifford algorithm for Boson Sampling [62]. Then, we count the number of collision-free outcomes over the 500 sampled outcomes and compute the average ratio of collision-free outcomes for each $N$ and circuit ensemble, where the average is taken over the $500$ sampled unitary circuits for each circuit ensemble.